150 W of power was used to move an object 22 meters with 15N of force. How long did it take ?
1
Expert's answer
2020-10-25T18:31:14-0400

The work done in this case is:


"A = Fs"

where "F = 15N" is acting force and "s = 22m" is the distance covered under this force. Thus, obtain:

By definition, the power is:


"P = \dfrac{A}{t}"

where "t" is the time of the process. Expressing "t", get:


"t = \dfrac{A}{P} = \dfrac{Fs}{P} = \dfrac{15N\cdot 22m}{150W} = 2.2s"

Answer. 2.2 s.